OpenStudy (anonymous):

What is the factorization of 3x^2 - 147?

OpenStudy (anonymous):

=3(x^2-49)=[difference of square a^2-b^2=(a-b)(a+b)] Now plug it in to see what you get. :)

OpenStudy (anonymous):

3(x-7)^2?

OpenStudy (owlcoffee):

Let's look at our ecuation: \[3x ^{2}-147\] We need to get rid of that 3 side to the x^2, so let's divide the whole ecuation by 3. We'll end up having: \[\frac{ 3x ^{2} }{ 3 }-\frac{ 147 }{ 3 }\]=\[x ^{2}-49\] With a little knowledge in factorization, we know that we are in front of a square difference: remember: \[a ^{2}-b ^{2}=(a+b)(a-b)\] Let's apply that. What we need to do is take the square root of the first and square root of the second and sum them, then do the same, but substracting them. \[(\sqrt{x ^{2}}+\sqrt{49})(\sqrt{x ^{2}}-\sqrt{49})\] So we'll end up having, the factorization: \[(x+7)(x-7)\] I hope it helped, good luck!

OpenStudy (anonymous):

I was way off lol

OpenStudy (anonymous):

No, you were close, because the correct form is 3 (x-7) (x+7)

OpenStudy (anonymous):

I understand now though. Thanks to the both of you. Much appreciation

OpenStudy (anonymous):

Your Welcome. :)

OpenStudy (owlcoffee):

forgot to multiply by 3 hehe <3

OpenStudy (anonymous):

Laughing out loud, :)